# Accurate Power Spectrum Estimation toward Nyquist limit

### v1

@@ 3 Numerical Tests
We conduct the numerical tests on the halo samples simulated by Gadget-4 code [10]. The simulation box length is $L=600\,h^{-1}\mathrm{Mpc}$. The cosmological parameters employed here are: $\Omega_{m}=0.308$, $\Omega_{\Lambda}=0.692$, $\Omega_{b}=0.049$, $H_{0}=67.8\,{\rm km}\,s^{-1}\,\rm{Mpc}^{-1}$. The halo sample is generated by the on-the-fly Friends-of-Friends halo finder. The sample consists of a total of 2,413,300 dark matter halos with mass ranging from $3.40\times 10^{11}M_{\odot}/h$ to $7.30\times 10^{14}M_{\odot}/h$, located at a redshift of $z=1$.
3.1 Full halo sample
@@
Figure 2: The power spectra of the full sample, with results for the CIC window function on the left and the TSC window function on the right. The top panel shows the power spectrum functions, while the middle and bottom panels depict the ratio of the power spectrum estimates to the results from the high-resolution grid. The blue lines represent the comparison between the two estimation methods without interlacing technique, and the red lines represent the comparison with interlacing technique. The dashed lines are calculated using the Nbodykit library, following the sequence of operations that involve correcting the window function first and then correcting the shot noise, according to the estimation formulae Eq. (2.10) and Eq. (2.17). Conversely, the solid lines follow Eq. (2.9) and Eq. (2.16). The red arrows in the graph indicate the frequency $k=\pi\sqrt[3]{n_{s}}$ that characterizes the sample’s average spatial separation. The gray region represents the statistical fluctuations of the power spectrum, i.e., ±$\pm\sigma=\pm\frac{P}{\sqrt{N_{\textrm{mode}}}}$, where $N_{\textrm{mode}}$ denotes the number of modes of each $k$ bin.
@@
Based on the aforementioned samples, we test the results using the CIC and TSC window functions with a grid size of $N_{g}=200$. The mesh size is $3\,h^{-1}{\rm Mpc}$ and the corresponding Nyquist frequency $k_{N}=1.047\,h{\rm Mpc}^{-1}$. We choose this relatively large mesh size under the consideration that for Stage-IV surveys a cubic box covering the whole footprint may reach $(6\,h^{-1}{\rm Gpc})^{3}$ and the analysis in performed on $2000^{3}$ mesh. We compare the differences among the four estimators mentioned above. To reflect the magnitude of systematic biases for the samples, we treat the results from a 10 times higher sampling rate, $N_{g}=2000$, as the true values for the power spectrum. We have confirmed that, at this high sampling frequency, the power spectrum results from different estimators with different window functions are overlapping with each other below the Nyquist frequency of $N_{g}=200$ mesh. Therefore, we can consider it as the true value for a low sampling frequency result. For the full sample with a total count of $N_{s}=2,413,300$ and a number density of $n_{s}=1.12\times 10^{-2}\,h^{3}{\rm Mpc}^{-3}$ , the power spectrum results are shown in Figure 2.
Under the CIC window, Eq. (2.9) yields an overestimated result, with an error exceeding 10% at $k=k_{N}$. Eq. (2.10) gives an underestimated result, with an error of around 10% near $k=k_{N}$. Both errors far exceed the level of statistical fluctuations. However, with the interlacing technique, Eq. (2.17) exhibits errors larger than statistical fluctuations only at the scale of $k=0.9k_{N}$, reaching around 1% at $k=k_{N}$. Meanwhile, Eq. (2.16) maintains an error within 0.1%.
@@
Under the TSC window, the errors are reduced. Eq. (2.9) provides an overestimated result, with an error of approximately 10% at $k=k_{N}$. Eq. (2.10) yields an underestimated result, with an error of around 5% near $k=k_{N}$. Both errors still significantly exceed the level of statistical fluctuations. However, with the interlacing technique, both Eq. (2.16) and Eq. (2.17) exhibit errors smaller than the level of statistical fluctuations within the Nyquist frequency. Specifically, Eq. (2.16) has an error within 0.002% at $k=k_{N}$, while Eq. (2.17) has an error within 0.1%.
We can observe that adopting higher-order window functions with the interlacing technique effectively reduces estimation errors, which is the main conclusion of [9]. However, when not using the interlacing technique, the reduction of errors with higher-order window functions is not significant. In terms of performance under the interlacing technique, Eq. (2.16) performs better than Eq. (2.17).
3.2 10% randomly selected halos
Figure 3: The power spectrum of the 10% randomly sampled data. The symbols have the same meaning as those in Figure 2.
We perform the same calculations as in Section 3.1, but this time we randomly selected 10% of the samples from the entire sample, resulting in a sample size of $N_{s}=241,330$. The sample number density is $n_{s}=1.12\times 10^{-3}\text{ Mpc}^{-3}h^{3}$. We obtain the power spectrum with the same amplitude as in Section 3.1, but the corresponding shot noise term $\frac{1}{N}$ is ten times larger after down-sampling, resulting in larger fluctuations on top of the original power spectrum. The results are shown in Figure 3.
Under the CIC window, Eq. (2.9) provides an overestimated result, and Eq. (2.10) yields an underestimated result. Both errors exceed 10% near $k=k_{N}$, far exceeding the level of statistical fluctuations. With the interlacing technique, Eq. (2.17) exhibits errors exceeding the level of statistical fluctuations below the scale of $k=0.7k_{N}$, reaching 10% at $k=k_{N}$, while Eq. (2.16) maintains an error within 1% despite larger fluctuations.
Under the TSC window, the errors are improved. Eq. (2.9) provides an overestimated result, with an error of approximately 10% at $k=k_{N}$. Eq. (2.10) yields an underestimated result, with an error of around 5% near $k=k_{N}$. Both errors still significantly exceed the level of statistical fluctuations. With the interlacing technique, Eq. (2.16) exhibits errors smaller than statistical fluctuations within the Nyquist frequency. Specifically, it has an error below 0.2% at $k=k_{N}$, while Eq. (2.17) exceeds the level of statistical fluctuations and reaches 0.7%.
Compared to the result in Section 3.1, the systematic bias is significantly larger due to the lower sample density, especially for the estimators with inexact shot noise subtraction. We can see that under the interlacing technique, the estimation Eq. (2.16) still performs better than Eq. (2.17).
3.3 10% high mass halos
@@
Figure 4: The power spectrum of a 10% subsample selected in high mass, with the same symbols as in Figure 2.
We perform the same calculations as in Section 3.1, but this time we select the top 10% of the samples in terms of mass from the entire sample, resulting in a sample size of $N_{s}=241,330$. The number density is $n_{s}=1.12\times 10^{-3}\text{ Mpc}^{-3}h^{3}$. High-mass halos tend to exhibit higher clustering than low-mass halos. Therefore, the power spectrum amplitudes are no longer the same as those obtained in Section 3.1. The large power spectrum amplitude of this sample makes the shot noise less important. However, the corresponding shot noise term $\frac{1}{N}$ unavoidably amplifies as the average number density decreases, resulting in larger fluctuations on top of the power spectrum. The results are shown in Figure 4.
Again, Eq. (2.9) provides an overestimated result, and Eq. (2.10) yields an underestimated result for CIC window. Both errors reach approximately 10% near $k=k_{N}$, far exceeding the level of statistical fluctuations. However, Eq. (2.9) still maintains an error of around 1% within $k=0.8k_{N}$, while Eq. (2.10) exceeds at larger scales. With the interlacing technique, Eq. (2.17) exhibits errors larger than statistical fluctuations below the scale of $k=0.7k_{N}$, reaching 9% at $k=k_{N}$. On the other hand, Eq. (2.16) fluctuates but remains within the level of statistical fluctuations. Under the TSC window, the errors are improved. Eq. (2.9) provides an overestimated result, with an error of approximately 7% at $k=k_{N}$. Eq. (2.10) yields an underestimated result, with an error of around 5% near $k=k_{N}$. Both errors still significantly exceed the level of statistical fluctuations. However, with the interlacing technique, Eq. (2.16) exhibits errors smaller than statistical fluctuations within the Nyquist frequency. Specifically, it has an error below 0.1% at $k=k_{N}$, while Eq. (2.17) exceeds statistical fluctuations only near the Nyquist frequency. We can see that under this sampling scheme, the estimation Eq. (2.16) still demonstrates superior accuracy.
3.4 1% randomly selected halos
@@
Figure 5: The power spectrum of the 1% randomly sampled data. The symbols have the same meaning as those in Figure 2.
We randomly select a sparse sample from the total dataset, consisting of 1% of the total sample size, which corresponds to $N_{s}=24,133$ halos. The number density of this sparse sample is $n_{s}=1.12\times 10^{-4}\,\text{Mpc}^{-3}h^{3}$. We perform the same computation as described in Section 3.1 on this sparse sample, and we expect to obtain power spectra with the same amplitude as in Section 3.1. However, due to the sparsity of this sample, the resulting power spectrum will exhibit even larger fluctuations compared to the results in Section 3.2. The results are presented in Figure 5. Under the CIC window, Eq. (2.10) overestimates the results, while Eq. (2.9) underestimates them. Both estimations show errors of around 10% near $k=0.7k_{N}$, which greatly exceeds the size of statistical fluctuations. Under the interlacing technique, Eq. (2.17) exhibits errors of over 50% at $k=0.9k_{N}$, while Eq. (2.16) exhibits errors of over 50% near the Nyquist frequency. Under the TSC window, Eq. (2.10) overestimates the results, with an error of over 50% at $k=0.95k_{N}$, and Eq. (2.9) underestimates them, with an error within 50% near the Nyquist frequency. Both estimations still greatly exceed the size of statistical fluctuations. Although under the interlacing technique the errors still exceed the size of statistical fluctuations. Eq. (2.16) exhibits errors within 5% near the Nyquist frequency, while Eq. (2.17) exhibits errors of around 15% near the same frequency. It can be seen that under this sampling scheme, the estimation using Eq. (2.16) still demonstrates a precision advantage.
3.5 1% high mass halos
@@
Figure 6: The power spectrum of a 1% subsample selected in high mass, with the same symbols as in Figure 2.
@@
The power spectrum of the 1% top mass halos, with a total sample size of $N_{s}=24,133$ and a number density of $n_{s}=1.12\times 10^{-4}\,\text{Mpc}^{-3}h^{3}$, has been computed using the same configuration as the previous sections. The results are presented in Figure 6. The power spectrum exhibits a higher level of clustering due to the selection of high-mass components. However, the estimated power spectrum is less than zero in the small scale even the fine grid is adopted. This nonphysical result is due to the fact that the halo distribution for this sample is sub-Poissonian. For example, in [4], for the most massive halo bin, the shot noise is only at a level of $2/3$ of the expected $1/N$ value. Thus, our esitmators naively over-subtract the shot noise term by assuming the shot noise being $1/N$. Despite the nonphysical reason for the negative power in the calculation, the comparison of different estimators is still valid in the sense that they are compared to the fine grid result. But the zero-crossing in the fiducial power spectrum leads to large fluctuations in the ratio plot. As a result, the values near that point are sensitive to how the frequency approaches it, leading to a false oscillation in the power spectrum ratio. We prefer not to compare the values near the zero-crossing scale. Overall, in the case of the 1% high mass data, the power spectrum estimation using the interlacing technique with Eq. (2.16) still provides the best results.

### v2

@@ 3 Numerical tests
We conduct the numerical tests on both the dark matter particle samples and the halo samples simulated by Gadget-4 code [10]. The simulation box length is $L=600\,h^{-1}\mathrm{Mpc}$. The cosmological parameters employed here are: $\Omega_{m}=0.308$, $\Omega_{\Lambda}=0.692$, $\Omega_{b}=0.049$, $H_{0}=67.8\,{\rm km}\,s^{-1}\,\rm{Mpc}^{-1}$. The dark matter sample contains totally $1024^{3}$ particles. The halo sample is generated by the on-the-fly Friends-of-Friends halo finder. A total of 2,413,300 dark matter halos are found with mass ranging from $3.40\times 10^{11}M_{\odot}/h$ to $7.30\times 10^{14}M_{\odot}/h$, located at a redshift of $z=1$.
3.1 Full dark matter sample
@@
Figure 2: The power spectra of the full dark matter sample, with results for the CIC window function on the left and the TSC window function on the right. All results extend up to the Nyquist frequency of its own. The top panel shows the power spectrum functions, while the middle and bottom panels depict the ratio of the power spectrum estimates to the results from the high-resolution grid. The blue lines represent the comparison between the two estimation methods without interlacing technique, and the red lines represent the comparison with interlacing technique. The dashed lines are calculated using the Nbodykit library, following the sequence of operations that involve correcting the window function first and then correcting the shot noise, according to the estimation formulae Eq. (2.10) and Eq. (2.17). Conversely, the solid lines follow Eq. (2.9) and Eq. (2.16). The gray region represents the statistical fluctuations of the power spectrum, i.e., ±$\pm\sigma=\pm\frac{P+1/n_{s}}{\sqrt{N_{\textrm{mode}}}}$, where $N_{\textrm{mode}}$ denotes the number of modes of each $k$ bin.
@@
The dark matter particle distribution in simulation is considered to be a Poissonian sampling of the underlying matter field. Thus we first test the results on the simulated dark matter samples using the CIC and TSC window functions with a grid size of $N_{g}=200$. The mesh size is $3\,h^{-1}{\rm Mpc}$ and the corresponding Nyquist frequency $k_{N}=1.047\,h{\rm Mpc}^{-1}$. We choose this relatively large mesh size under the consideration that for Stage-IV surveys a cubic box covering the whole footprint may reach $(6\,h^{-1}{\rm Gpc})^{3}$ and the analysis will be performed on $2000^{3}$ mesh. We compare the differences among the four estimators mentioned above. To reflect the magnitude of systematic biases for the samples, we treat the results from a 10 times higher sampling rate, $N_{g}=2000$, as the true values for the power spectrum. We have confirmed that, at this high sampling frequency, the power spectrum results from different estimators with different window functions are overlapping with each other below the Nyquist frequency of $N_{g}=200$ mesh. Therefore, we can consider it as the true value for a low sampling frequency result.
For the full sample with a total count of $N_{s}=1,073,741,824$ and a number density of $n_{s}=4.97\,h^{3}{\rm Mpc}^{-3}$ , the power spectrum results are shown in Figure 2. Under the CIC window, Eq. (2.9) yields an overestimated result, with an error exceeding 10% at $k=0.9k_{N}$. Eq. (2.10) gives an underestimated result, with an error of around 6% near $k=k_{N}$. The statistical uncertainty of the band power is given by $\sigma=\frac{P+1/n_{s}}{\sqrt{N_{\textrm{mode}}}}$, where $N_{\textrm{mode}}$ denotes the number of modes in each $k$ bin. Note that although we are attempting to clearly subtract the mean shot noise from the measurement, the shot noise residual still contribute to the statistical uncertainty. Both errors far exceed the level of statistical uncertainty level denoted by grey band. Note that the difference between the two estimators fully depends on the different corrections applied to the signal term, since the shot noise level under such circumstances, $\frac{1}{n_{s}}=0.20\ h^{-3}{\rm Mpc^{3}}$, is insignificant compared to the signal. Yet with the interlacing technique, both estimators Eq. (2.17) and Eq. (2.16) remain within the statistical fluctuations at the Nyquist frequency. The overlapping between the two is expected since Eq. (2.17) and Eq. (2.16) is exactly same when shot noise is fully negligible. The small deviation ($\sim 0.25\%$ at $k=k_{N}$) from unity is the result of the aliasing effect on the signal, even the interlacing technique is adopted. In Appendix B we provide high-order interlacing schemes which can further reduce the deviation.
@@
Under the TSC window, the systematic biases are reduced. Eq. (2.9) provides an overestimated result, with an error of approximately 18% at $k=k_{N}$. Eq. (2.10) yields an underestimated result, with an error of 3% near $k=k_{N}$. Both errors still significantly exceed the level of statistical fluctuations. However, with the interlacing technique, both Eq. (2.16) and Eq. (2.17) exhibit errors much smaller than the level of statistical fluctuations within the Nyquist frequency, and also far below $1\%$.
3.2 Down-sampled dark matter
@@
Figure 3: The power spectrum of the dark matter sampled down to $n_{s}=1.12\times 10^{-2}\,h^{3}{\rm Mpc}^{-3}$. The red arrows in the graph indicate the frequency $k=\pi\sqrt[3]{n_{s}}$ that characterizes the sample’s average spatial separation, and the green arrows indicate the scale where shot noise level surpass that of the signals. Rest of the symbols have the same meaning as those in Figure 2. With interlacing, the estimator Eq. (2.16) (red solid line) has much better performance than Eq. (2.17) (red dashed line) for CIC window.
To highlight the difference among the various estimators towards low number density limit, we randomly select three subsamples of $N_{s}=2,413,300$, $N_{s}=241,330$, and $N_{s}=24,133$ from the entire dark matter data set. The sample size can be kind of arbitrary but we specifically chose the above sample sizes for the comparison with later-discussed halo samples of the same sizes. For the sample with a total count of $N_{s}=2,413,300$ and a number density of $n_{s}=1.12\times 10^{-2}\,h^{3}{\rm Mpc}^{-3}$, the power spectrum results are shown in Figure 3.
For both CIC and TSC window without interlacing, Eq. (2.9) and Eq. (2.10) has similar result as the full sample. The reason is that the number density here is still sufficiently high, and thus the main problem is the aliasing in the signal term. When interlacing is adopted, the performance of shot noise subtraction becomes important. Under the CIC window, Eq. (2.17) exhibits errors larger than statistical fluctuations at the scale of $k>0.9k_{N}$, reaching around 1% at $k=k_{N}$. Meanwhile, Eq. (2.16) maintains an error within 0.2%, and below the statistical uncertainty. While for TSC window with interlacing, the bias from the two estimators is almost negligible.
For this number density, we can observe that adopting higher-order window functions with the interlacing technique effectively reduces estimation errors, which is the main conclusion of [9]. However, when not using the interlacing technique, the reduction of errors with higher-order window functions is significant but not sufficiently good. In terms of performance under the interlacing technique, Eq. (2.16) performs better than Eq. (2.17).
@@
Figure 4: The power spectrum of the dark matter sampled down to $n_{s}=1.12\times 10^{-3}\text{ Mpc}^{-3}h^{3}$. The symbols have the same meaning as those in Figure 3. With interlacing, the estimator Eq. (2.16) (red solid line) begin to show better result than Eq. (2.17) (red dashed line) even for TSC window.
As for the sample with a total count of $N_{s}=241,330$ and a number density of $n_{s}=1.12\times 10^{-3}\text{ Mpc}^{-3}h^{3}$, we obtain the power spectrum with the same amplitude as in Figure  3, but the corresponding shot noise term $\frac{1}{N}$ is ten times larger after down-sampling, resulting in larger fluctuations on top of the original power spectrum. The results are shown in Figure 4.
Under the CIC window, Eq. (2.9) provides an overestimated result, with an error reaching 30% at $k=k_{N}$. Eq. (2.10) yields an underestimated result, with an error reach 10% at $k=k_{N}$. Both errors far exceed the level of statistical fluctuations. The increased systematic bias compared to the result in Figure 3 is fully due to the worse shot noise subtraction induced by the estimators, since the underlying aliasing contaminated signal term is the same. With the interlacing technique, Eq. (2.17) exhibits errors exceeding the level of statistical fluctuations around the scale of $k=0.8k_{N}$, reaching 10% at $k=k_{N}$, while Eq. (2.16) maintains an error within 0.6% despite some fluctuations. Again, this also explicitly means that Eq. (2.17) is far from optimal in terms of shot noise subtraction.
Under the TSC window, the errors are improved. Eq. (2.9) provides an overestimated result, with an error of approximately 18% at $k=k_{N}$. Eq. (2.10) yields an underestimated result, with an error of around 5% near $k=k_{N}$. Both errors still significantly exceed the level of statistical fluctuations. With the interlacing technique, Eq. (2.16) exhibits errors smaller than statistical fluctuations within the Nyquist frequency. Specifically, it has an error below 0.2% at $k=k_{N}$. Meanwhile, the result from Eq. (2.17) is also below the statistical uncertainty level, but reach 0.6% at $k=k_{N}$. We emphasize that $n_{s}=1.12\times 10^{-3}\text{ Mpc}^{-3}h^{3}$ is the number density that estimator Eq. (2.16) show obvious better performance than Eq. (2.17) ($0.2\%$ bias versus $0.6\%$) with TSC window and interlacing technique.
@@
Figure 5: The power spectrum of the dark matter sampled down to $n_{s}=1.12\times 10^{-4}\,\text{Mpc}^{-3}h^{3}$. The symbols have the same meaning as those in Figure 3.
As for the sparse sample with a total count of $N_{s}=24,133$ and a low number density of $n_{s}=1.12\times 10^{-4}\,\text{Mpc}^{-3}h^{3}$, we perform the same computation as described above, and we expect to obtain power spectra with the same amplitude as the former two samples. However, due to the sparsity of this sample, the resulting power spectrum will exhibit even larger fluctuations compared to the results in Figure 4. The results are presented in Figure 5. Under the CIC window, Both Eq. (2.10) and Eq. (2.9) show errors around 10% for $k>0.65k_{N}$, which is only half of the statistical fluctuation level. Under the interlacing technique, Eq. (2.17) overestimates the results, exhibits an error of over 50% at $k=0.9k_{N}$, while Eq. (2.16) underestimates it, exhibits an error below 10%. Careful readers may ask why the estimator Eq. (2.17) with interlacing technique performs much worse than the estimators Eq. (2.10) and Eq. (2.9) without interlacing. The reason is that Eq. (2.10) and Eq. (2.9) correctly take into account the aliasing on the shot noise and subtract it (i.e. $\frac{1}{N}C({\boldsymbol{k}})$) from the measurement. On the contrary, Eq. (2.17) treats the shot noise to be $\frac{1}{N}W^{2}({\boldsymbol{k}})$ and ignore the aliasing effect on the shot noise. Thus, the shot noise is greatly underestimated and the residual is large even though the interlacing is adopted. This example demonstrates that for the shot noise dominated case, the correct subtraction of shot noise is vital.
Under the TSC window, both Eq. (2.10) and Eq. (2.9) give estimations within the size of statistical fluctuations. However, the fluctuations in the ratio are still obvious. Under the interlacing technique, again the two estimators have errors smaller than the size of statistical fluctuations, but the above fluctuations in the ratio are less obvious. Thanks to the large statistical noise level of $25\%$, it seems to be safe to use both estimators. Note that in terms of the absolute value of the systematic bias, Eq. (2.16) exhibits errors within 2% near the Nyquist frequency, while Eq. (2.17) exhibits errors of around 8% near the same frequency. It can be seen that under this low sampling rate, the estimation using Eq. (2.16) still demonstrates a precision advantage.
3.3 Halo sample
Refer to caption Refer to caption
Figure 6: The power spectrum of a high-mass halo subsample ($n_{s}=1.12\times 10^{-3}\text{ Mpc}^{-3}h^{3}$), with the same symbols as in Figure 3. For all the estimators, the deviation is slightly smaller than the corresponding dark matter sample, thanks to boosted power amplitude by the halo bias.
In this section we perform the same calculations as in Section 3.1 & 3.2, but on the halo samples. We do not repeat the random down-sampling process on the full halo sample as what we did in Section 3.2 since the behavior has been well tested there. Instead, we gradually increase the mass threshold of the haloes to decrease the sample size. Note that it has been pointed out in the literature (for example, see [Baldauf2013] and [4]) that the shot noise contribution in halo power spectrum is in general not Poissonian, particularly for high-mass haloes. As a consequence, the large scale power is lower than the expected $1/n_{s}$ and gradually transits to the Poisson expectation in the small-scale. This means that at least for large scales, the shot noise is over-subtracted in the four estimators we presented. Here we focus on small scales and this part serves as a performance test on the competition effects between the halo bias which boosts the signal and the shot noise which contaminate the signal.
We first select the top 10% of the samples in terms of mass from the entire halo sample, resulting in a sample size of $N_{s}=241,330$. The number density is same as the dark matter sample in Figure 4, $n_{s}=1.12\times 10^{-3}\text{ Mpc}^{-3}h^{3}$. High-mass halos tend to exhibit higher clustering than dark matter. Therefore, the power spectrum amplitude is boosted. The large power spectrum amplitude of this sample makes the shot noise less important than the dark matter sample with the same sample size. The results are shown in Figure 6.
Again, Eq. (2.9) provides an overestimated result, and Eq. (2.10) yields an underestimated result for CIC window. Both errors reach approximately 10% near $k=k_{N}$, far exceeding the level of statistical fluctuations. However, Eq. (2.9) still maintains an error of around 1% within $k=0.8k_{N}$, while Eq. (2.10) exceeds at larger scales. Compared to the results in Figure 4, the systematic deviation is indeed slightly reduced thanks to the halo bias of the sample. With the interlacing technique, Eq. (2.17) exhibits errors larger than statistical fluctuations below the scale of $k=0.85k_{N}$, reaching 9% at $k=k_{N}$. On the other hand, Eq. (2.16) fluctuates but remains within the level of statistical fluctuations. Under the TSC window, the errors are improved. Eq. (2.9) provides an overestimated result, with an error of approximately 7% at $k=k_{N}$. Eq. (2.10) yields an underestimated result, with an error of around 5% near $k=k_{N}$. Both errors still significantly exceed the level of statistical fluctuations. With interlacing, both estimators, Eq. (2.16) and Eq. (2.17), present deviations far less than the statistical uncertainty level. Careful readers can find that around the Nyquist frequency the estimation Eq. (2.16) still demonstrates superior accuracy.
Refer to caption Refer to caption
@@ 3.3 Halo sample
Figure 7: The power spectrum of a $n_{s}=1.12\times 10^{-4}\,\text{Mpc}^{-3}h^{3}$ subsample selected in high halo mass, with the same symbols as in Figure 3.
@@
The power spectrum result for the top 1% massive halos, with a total sample size of $N_{s}=24,133$ and a number density of $n_{s}=1.12\times 10^{-4}\,\text{Mpc}^{-3}h^{3}$, is presented in Figure 7. The power spectrum exhibits a higher level of clustering due to the selection of high-mass components. However, the estimated power spectrum is less than zero in the small scale even the fine grid is adopted. This counterintuitive result is due to the fact that the halo distribution for this sample is sub-Poissonian. For example, in [4], for the most massive halo bin, the shot noise is only at a level of $2/3$ of the expected $1/N$ value. Thus, our estimators naively over-subtract the shot noise term by assuming the shot noise being $1/N$. Despite the negative power in the calculation, the comparison of different estimators is still valid in the sense that they are compared to the fine grid result. But the zero-crossing in the fiducial power spectrum leads to large fluctuations in the ratio plot. As a result, the values near that point are sensitive to how the frequency approaches it, leading to a false oscillation in the power spectrum ratio. We prefer not to compare the values near the zero-crossing scale. Overall, in the case of the 1% high mass data, the power spectrum estimation using the interlacing technique with Eq. (2.16) still provides the best results.